OpenStudy (anonymous):

Determine whether the lines through the pairs of points are parallel, perpendicular, or neither. 3x - 2y = -5 2x + 3y = -15 Choose one answer. a. parallel b. neither c. perpendicular

OpenStudy (unklerhaukus):

3x - 2y = -5 \[\rightarrow y=(3/2)x +5/2\] 2x + 3y = -15 \[\rightarrow y=(-2/3)x-5\] the gradients are \[m_1={-1 \over m_2}\] which means they lines are perpendicular